OpenStudy (anonymous):

Can you please tell me how to set this up? Thanks! The Pretty Pizza Company wants to open a new pizza place either on the north side of town or the south side of town. They want to open the pizza place where the ratio of pizza places per person is currently less than 0.008. The population of 15 blocks of the north side of town is 220 people per block, but there are already 26 pizza places in the area. The population of 22 blocks of the south area is 175 people per block, and there are 33 pizza places in the area. Where, if anywhere, should the Pretty Pizza Company open a new pizza place?

OpenStudy (phi):

ratio of pizza places per person means write the fraction \[ \frac{\text{# of pizza places}}{\text{person} } \] can you do that for both sides of town ?

OpenStudy (phi):

For example, on the north side there are 26 pizza places. that goes in the top of the fraction For the bottom, the # of people, we need the # of people in the north side The north side has 15 blocks with 220 people per block,

OpenStudy (anonymous):

I did that, and the answer for each is over the percentage given, however that answer choice is incorrect (neither side of town).

OpenStudy (phi):

for the north side, you should get \[ \frac{26}{15\cdot 220} = 0.00788 \]

OpenStudy (phi):

for the south side 33/(175*33) = 0.0057

OpenStudy (anonymous):

Oh, I just did the people not the blocks, I thought that was extraneous! Thank you so much!

OpenStudy (phi):

although if I use the correct numbers: in the south 22 blocks of the south area is 175 people per block, 22 blocks * 175 = 3850 and 33 pizza places in the area gives 33/3850 = 0.00857 So what we have is North side ratio is 0.00788 South side ratio is 0.00857 and it's ok to open if the ratio is *currently* less than 0.008. that means the north side is ok.

OpenStudy (anonymous):

UGH, just submitted that response and it is wrong. Not sure why, because it makes sense, unless they are rounding up on north side.